HETAV PATEL

Grade 11,

Fill in the boxes: ▯ - 1 / 5 = 1 / 2 Fill in the boxes: ▯ - 1 / 5 = 1 / 2

 Fill in the boxes: 
 ▯ - 1 / 5 = 1 / 2
Fill in the boxes: 
 ▯ - 1 / 5 = 1 / 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

▯- 1 / 5 = 1 / 2
▯=1 / 2 + 1 / 5
▯= [(1 × 5) + (1 × 2)] / 10
▯= (5 + 2) / 10
▯= 7 / 10

​Thanks